Fermi National Laboratory
 

Accelerator Update for July 5, 2006 - July 6, 2006

 

Wednesday July 5

The day shift began with Operations monitoring store 4813 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam store, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with NuMI off due to a horn cooling problem, and with MiniBooNE taking beam.

Operations established beam to the Meson MTest experiment at 10:07 AM.

Operations turned off all beam at 11:48 AM and began preparing the Preaccelerator , Linac, Booster, and MI for accesses.

The access into the H- Source ended at 1:21 PM.

Linac came out of access at 2:27 PM.

Booster came out of access at 4:56 PM.

MiniBooNE began taking beam at 5:35 PM.

Pbar began stacking at 5:39 PM.

At 5:43 PM, Operations reported that two cold compressors had tripped off due to over heating for the TeV sector A0 area. Cryo system technicians turned on a compressor at Switchyard (SY) and two other compressors (at D0 and E0) to compensate. The Duty Mechanic reported that this was a flow issue due to a clogged pond strainer.

At 6:34 PM, Operations halted beam for fifteen minutes so that techs could investigate a problem with the Booster ORBUMP power supply.

Pbar resumed stacking at 6:49 PM.

At 7:15 PM, Operations reported that the SY work wasn’t going well. Operations decided to prepare the Recycler to stash antiprotons.

The stash ended and Pbar resumed at 8:13 PM.

A FESS expert reported at 8:37 PM that they had added a fire hose at TeV sector A0 to improve water pressure. However, while doing this, the water pressure returned to normal; the clog released on its own.

Thursday July 6

The midnight shift began with Operations monitoring store 4813 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler began preparing to stash at 12:31 AM. The stash ended. Pbar resumed stacking at 1:13 AM.

At 1:17 AM, all of the quadrupole power supplies associated with the first Linac RF station (LRF1) alarmed. Operators replaced a Timing Pulse Repeater module.

At 4:32 AM, a Linac RF station (KRF6) tripped off on LCW flow. Operators were able to persuade the flow meter to resume normal operations.

The Recycler began preparing to stash at 5:17 AM. The stash ended. Pbar resumed stacking at 5:47 AM.

The Plans for Thursday

The plans for today are to terminate store 4813 at 1 PM and then allow TeV experts to conduct a two-hour study; a store will follow the study.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/06/2006    email Fermilab